OpenStudy (lilday12):

For each equation below, determine whether the number shown in bold is a solution to the equation. If the number is not a solution, then find the solution. Show your work. A. x + 17 = -3; -20 B. y - 12 = 14; 10 Part I: Plug each answer into the equation for the variable. (2 points; 1 point each)

OpenStudy (kkutie7):

well for A you should start by plugging in the "bold" number in for the x

OpenStudy (lilday12):

-20+17=-3

OpenStudy (kkutie7):

here: \[A: (-20)+17=-3\] \[B:(10)-12=14\] solve the left side and see if it matches the right.

OpenStudy (kkutie7):

are you there?

OpenStudy (lilday12):

Yes

OpenStudy (kkutie7):

let me see some work

OpenStudy (lilday12):

Determine if the number given is a correct solution. If the number is not a solution, find the correct solution.